# KCL-PH-TH/2025-19

### v2

@@ 1 Introduction
Sub-GeV DM candidates, with masses in the MeV–GeV range, are theoretically well-motivated and arise in a variety of extensions of the Standard Model [7, 12, 13]. However, their low mass poses detection challenges. Direct detection experiments lose sensitivity in this regime, while indirect searches are hindered by charged particle suppression in the heliosphere and the limited sensitivity of existing $\gamma$-ray instruments in the keV–GeV band, the so-called “MeV gap” [14, 12, 15]. Nevertheless, DM annihilation or decay into $e^{\pm}$ can inject low-energy electrons and positrons into the interstellar medium. These secondaries interact with Galactic radiation fields via bremsstrahlung radiation and inverse Compton (IC) scattering, producing X-rays in the keV to MeV range [1], which offers a valuable observational window.
@@
In this work, we address this gap by placing new constraints on light DM and sub-asteroid mass PBHs using eROSITA’s full-sky X-ray maps. We model continuum emission from DM-induced low-energy $e^{\pm}$ and Hawking evaporation from PBHs, deriving upper limits on the annihilation cross section, decay lifetime, and PBH DM fraction across a wide mass range. We then compare these constraints with existing bounds from other probes, demonstrating the complementarity and reach afforded by eROSITA. Additionally, we revisit constraints in Refs. [1, 2] derived from XMM-Newton data [28], showing that applying the geometric solid angle instead of the exposure-weighted solid angle led to flux differences of approximately four orders of magnitude. After addressing this, the XMM-Newton limits are found to be weaker than those from eROSITA, realigning the landscape of X-ray DM constraints. We note that the above discussion above does not apply to Ref. [28] which derives constraints using line-like signals from decaying DM from XMM-Newton.
@@
The maps are produced in HEALPix format, using the healpy package [30], with a pixel size of $0.22\degree\times 0.22\degree$ in units cts/s/$\mathrm{deg}^{2}$ (and subsequently converted to keV/sr/$\mathrm{cm^{2}}$/s for the purposes of the analysis) after properly filtering out point sources (according to the eRASS1 source catalog [31]) above a selected flux threshold (point sources with ML_FLUX_P2+ML_FL UX_P3 in the eRASS1 source catalog above $10^{-13}$ were removed) and removing the particle background. This threshold is selected such that 19% of the cosmic X-ray background is removed, as discussed in Fig. 12 of the main data release collaboration paper [31]. Overall, the individual steps outlined by the eROSITA collaboration in Section 2 of Ref. [32] were adapted to obtain the corresponding diffuse X-ray map from the selected region.
@@ 3.1 Sub-GeV dark matter annihilation and decay
We restrict our analysis to the channel of direct $e^{+}e^{-}$ production, from a minimum kinematically allowable mass of $m_{e}$ ($2m_{e}$) for annihilation (decay) respectively. In both cases, we take the maximum mass of DM to be $1000$ MeV. We use a 2D axisymmetric Galactic model with a NFW DM profile [36] representing our base case. Other decay channels including the production of pions and muons can also be considered, but the constraints become much weaker than in previous analyses  [1, 2] for the high DM masses needed. The propagation parameters used in this work are obtained from a fit to AMS-02 data [37, 38], and are listed in Table I of Ref. [2], where more of the technical details are given. The interplay between diffusion and energy losses is key at sub-GeV energies. Energy losses in the DRAGON2 code include ionisation, Coulomb, bremsstrahlung, IC, synchrotron and in-flight positron annihilation processes. The spatial and energy resolution employed is $\sim 150$ pc and $5$%.
The resulting $e^{\pm}$ distributions are passed to the HERMES code [39] to compute IC X-ray emission, integrating over the ISRFs [40]. Although IC losses have little effect on the $e^{\pm}$ spectra, the emission dominates the X-ray signal, especially below the MeV range that is relevant for experiments such as eROSITA and XMM-Newton. We estimate the ISRF modelling uncertainty to be below 30% [41], and we separately account for internal bremsstrahlung, also known as final-state radiation (FSR), as in Ref. [1]. Full simulation inputs and outputs are available in Ref. [42].
@@ 3.3 Dark matter density profiles
where $\rho_{s}$ is the characteristic density, set to from the assumption that the local DM density is of $0.4$ GeV/$\mathrm{cm^{3}}$ and $r_{s}=20~{}\text{kpc}$ [53].
@@
where the normalisation $\rho_{s}$ is again given by fixing the local DM density to $0.4$ GeV/$\mathrm{cm^{3}}$, the scale radius is the same as for the NFW case and we set the inner slope to $\gamma=1.4$.
@@ 4 Constraints from eROSITA data
[…]
@@
In Fig. 3, we show the constraints on the DM annihilation cross section, $\langle\sigma v\rangle$, and decay lifetime, $\tau$, into electrons and positrons, assuming a nominal NFW DM profile, using our combined ring analysis (orange line). The orange uncertainty bands represent those arising from varying of the propagation parameters in the scenarios described above. We also show existing constraints from INTEGRAL, NuSTAR and Suzaku (green lines) [1, 56, 57]. Compared to these probes, eROSITA provides the strongest constraints for DM masses up to a few tens of MeV. If we take constraints from rings individually, we find that rings 1 - 3 provides the strongest constraints, nearly identical to constraints derived from the full ring analysis for annihilation, as opposed to decay which benefits more from using all rings. As we observe, the uncertainties are larger for lower DM masses, which is due to the impact of reacceleration in the $e^{\pm}$ products: the lower their energy, the more prone they are to being boosted by the interaction with plasma waves [58, 59], as extensively reported and discussed in Refs. [2, 15].
To highlight the importance of these constraints, we show a comparison of the decay constraints obtained with eROSITA in comparison with those obtained from cosmological measurements, heating of dwarf galaxies and direct observations of CRs (in particular, from Voyager-1) and diffuse X- and $\gamma$-rays. We are not showing constraints from the 511 keV line [60] in this plot, which can reach lifetimes of $10^{29}$ s for $1$ MeV DM particles. The cosmological constraints are from CMB measurements and are obtained from the study of anisotropies in the CMB power spectrum [61, 62], the constraint from Leo-T dwarf galaxy was derived requiring the rate of heat injection by DM to not exceed the ultra-low radiative cooling rate of gas observed [63], and those from diffuse $\gamma$-rays are constraints from the FSR emission associated with DM [64] (obtained with COMPTEL, INTEGRAL and HEAO-I). Meanwhile, the constraint from Voyager-1 data is obtained by comparing the expected interstellar flux of electrons and positrons from DM with the flux measurements on board the probe. For the latter, it is important to note that in conservative conditions where reacceleration is very low, no limit can be obtained for DM masses below $\sim 20$ MeV. As we see, the eROSITA constraints are especially competitive below $10$ MeV, although at higher masses they are weaker. However, we emphasise that these limits are quite conservative, given that we directly mask the inner ±$\pm 5^{\circ}$ around the Galactic plane [^1] When doing the analysis adopting a ±$\pm 2^{\circ}$ mask, our constraints improve by no more than a factor of 2.. On top of this, we note that including the astrophysical background components could lead to up to even an order of magnitude stronger constraints from eROSITA.
@@ 5 Revising XMM-Newton constraints
[…]
@@
To emphasise this, in Fig. 5 we plot the ring 3 fluxes (using a comparable $|b|<2^{\circ}$ inner mask of the Galactic plane) recorded for both eROSITA (orange points) and XMM-Newton assuming the exposure-weighted average solid angle (black points) as intended in the data analysis and the geometric solid angle (grey points) in the energy range 00-$6$ keV. We use experiment independent units of $\mathrm{keV\ s^{-1}\ sr^{-1}\ cm^{-2}}$ to produce compatible spectra. We see that using the geometric solid angle significantly underestimates the true flux and, in fact, XMM-Newton should provide somewhat weaker or comparable constraints to eROSITA once this has been addressed.
@@
For annihilation, the revised XMM-Newton bounds are roughly an order of magnitude weaker than those from eROSITA across the full DM mass range. For decaying DM, XMM-Newton constraints are more competitive at low masses, where FSR dominates, but become nearly two orders of magnitude weaker at higher masses, where IC emission governs the size of the signal. For PBHs, we find that XMM-Newton constraints are consistently about an order of magnitude weaker than those from eROSITA over the full PBH mass range considered. This clarification realigns the role of XMM-Newton in the broader landscape of X-ray DM searches and confirms eROSITA as the current leading X-ray instrument for probing sub-GeV dark matter and evaporating PBHs.
@@ 6 Discussion and conclusions
We find that, even with just six months of data, eROSITA already provides X-ray constraints among the most stringent to date. For decaying DM with mass below $\sim 10$ MeV, our limits surpass those from CMB and diffuse $\gamma$-ray observations. However, in this low-mass regime, signal predictions are subject to larger uncertainties due to charged particle propagation modeling. Looking ahead, eROSITA’s ability to resolve point sources will improve with future data releases, allowing for better subtraction of unresolved backgrounds and potentially yielding significantly stronger constraints. Current diffuse emission is dominated by unresolved sources such as coronal stellar emission and accreting white dwarfs. Improved estimates of this unresolved component could tighten the bounds presented here by orders of magnitude.

### v3

@@ 1 Introduction
Sub-GeV DM candidates, with masses in the MeV–GeV range, are theoretically well-motivated and arise in a variety of extensions of the Standard Model [53, 17, 19, 1]. However, their low mass poses detection challenges. Direct detection experiments lose sensitivity in this regime, while indirect searches are hindered by charged particle suppression in the heliosphere and the limited sensitivity of existing $\gamma$-ray instruments in the keV–GeV band, the so-called “MeV gap” [22, 17, 12]. Nevertheless, DM annihilation or decay into $e^{\pm}$ can inject low-energy electrons and positrons into the interstellar medium. These secondaries interact with Galactic radiation fields via bremsstrahlung radiation and inverse Compton (IC) scattering, producing X-rays in the keV to MeV range [18], which offers a valuable observational window.
@@
In this work, we address this gap by placing new constraints on light DM and sub-asteroid mass PBHs using eROSITA’s full-sky X-ray maps. We model continuum emission from DM-induced low-energy $e^{\pm}$ and Hawking evaporation from PBHs, deriving upper limits on the annihilation cross section, decay lifetime, and PBH DM fraction across a wide mass range. We then compare these constraints with existing bounds from other probes, demonstrating the complementarity and reach afforded by eROSITA. Additionally, we revisit constraints in Refs. [18, 23] derived from XMM-Newton data [38], showing that applying the geometric solid angle instead of the exposure-weighted solid angle led to flux differences of approximately four orders of magnitude. After addressing this, the XMM-Newton limits are found to be weaker than those from eROSITA, realigning the landscape of X-ray DM constraints. We note that the above discussion does not apply to Ref. [38] which derives constraints using line-like signals from decaying DM from XMM-Newton.
@@
[…]
@@ 3.1 Sub-GeV dark matter annihilation and decay
We restrict our analysis to the channel of direct $e^{+}e^{-}$ production, from a minimum kinematically allowable mass of $m_{e}$ ($2m_{e}$) for annihilation (decay) respectively. In both cases, we take the maximum mass of DM to be $1000$ MeV. We use a 2D axisymmetric Galactic model with a NFW DM profile [50] representing our base case. Other annihilation (decay) channels such as $\mu^{+}\mu^{-}$ and $\pi^{+}\pi^{-}$ can also be considered. While these yield softer secondary $e^{\pm}$ spectra than direct $e^{+}e^{-}$, their production thresholds require higher DM masses. The resulting $e^{\pm}$ are therefore injected at higher absolute energies, in a regime where constraints from instruments such as INTEGRAL already dominate, making the eROSITA bounds comparatively weaker [18, 23]. The propagation parameters used in this work are obtained from a fit to AMS-02 data [27, 47], and are listed in Table I of Ref. [23], where more of the technical details are given. The interplay between diffusion and energy losses is key at sub-GeV energies. Energy losses in the DRAGON2 code include ionisation, Coulomb, bremsstrahlung, IC, synchrotron and in-flight positron annihilation processes. The spatial and energy resolution employed is $\sim 150$ pc and 55%.
The resulting $e^{\pm}$ distributions are passed to the HERMES code [31] to compute IC X-ray emission, integrating over the ISRFs [59].
Although IC losses have little effect on the $e^{\pm}$ spectra, the emission dominates the X-ray signal, especially below the MeV range that is relevant for experiments such as eROSITA and XMM-Newton. We estimate the ISRF modelling uncertainty to be below 30% [60], and we separately account for final-state radiation (FSR) from the outgoing $e^{\pm}$, see Eq. (6) of Ref. [17]. This contribution is sometimes referred to as “internal bremsstrahlung” in the DM literature (eg. Ref. [13]). We emphasise that we do not include additional model-dependent internal bremsstrahlung from intermediate states, which depends on the underlying particle physics model. Full simulation inputs and outputs are available in Ref. [28].
@@ 3.3 Dark matter density profiles
[…]
@@
[…]
@@ 4 Constraints from eROSITA data
$\chi^{2}=\sum_{i}\left(\frac{\mathrm{Max}[\phi_{\mathrm{DM},i}(p,m_{\chi})-\phi_{\mathrm{eROSITA},i},\ 0]}{\sigma_{\mathrm{eROSITA},i}}\right)^{2}$ (4.1)
@@
In Fig. 3, we show the constraints on the DM annihilation cross section, $\langle\sigma v\rangle$, and decay lifetime, $\tau$, into electrons and positrons, assuming a nominal NFW DM profile, using our combined ring analysis (orange line). The orange uncertainty bands represent those arising from varying of the propagation parameters in the scenarios described above. We also show existing constraints from INTEGRAL, NuSTAR and Suzaku (green lines) [18]. Compared to these probes, eROSITA provides the strongest constraints for DM masses up to a few tens of MeV. If we take constraints from rings individually, we find that rings 1 - 3 provide the strongest constraints, nearly identical to constraints derived from the full ring analysis for annihilation, as opposed to decay which benefits more from using all rings. As we observe, the uncertainties are larger for lower DM masses, which is due to the impact of reacceleration in the $e^{\pm}$ products: the lower their energy, the more prone they are to being boosted by the interaction with plasma waves [57, 55], as extensively reported and discussed in Refs. [23, 12].
To highlight the importance of these constraints, we show a comparison of the decay constraints obtained with eROSITA in comparison with those obtained from cosmological measurements, heating of dwarf galaxies and direct observations of CRs (in particular, from Voyager-1) and diffuse X- and $\gamma$-rays. We are not showing constraints from the 511 keV line [24] in this plot, which can reach lifetimes of $10^{29}$ s for $1$ MeV DM particles. The cosmological constraints are from CMB measurements and are obtained from the study of anisotropies in the CMB power spectrum [56, 46], the constraint from Leo-T dwarf galaxy was derived requiring the rate of heat injection by DM to not exceed the ultra-low radiative cooling rate of gas observed [61], and those from diffuse $\gamma$-rays are constraints from the FSR emission associated with DM [33] (obtained with COMPTEL, INTEGRAL and HEAO-I). Meanwhile, the constraint from Voyager-1 data is obtained by comparing the expected interstellar flux of electrons and positrons from DM with the flux measurements on board the probe. For the latter, it is important to note that in conservative conditions where reacceleration is very low, no limit can be obtained for DM masses below $\sim 20$ MeV. As we see, the eROSITA constraints are especially competitive below $10$ MeV, although at higher masses they are weaker. However, we emphasise that these limits are quite conservative, given that we directly mask the inner $\pm 5^{\circ}$ around the Galactic plane [^1] When doing the analysis adopting a $\pm 2^{\circ}$ mask, our constraints improve by no more than a factor of 2.. On top of this, we note that including the astrophysical background components could lead to up to even an order of magnitude stronger constraints from eROSITA.
@@ 5 Revising XMM-Newton constraints
$\Omega_{\mathrm{exp}}=\int_{\textrm{R}}d\Omega\frac{T_{\mathrm{eff}}(\Omega)}{T_{\mathrm{tot}}(\Omega)},$ (5.1)
@@
To emphasise this, in Fig. 5 we plot the ring 3 fluxes (using a comparable $|b|<2^{\circ}$ inner mask of the Galactic plane) recorded for both eROSITA (orange points) and XMM-Newton assuming the exposure-weighted average solid angle (black points) as intended in the data analysis and the geometric solid angle (grey points) in the energy range 0-66 keV. We use experiment independent units of $\mathrm{keV\ s^{-1}\ sr^{-1}\ cm^{-2}}$ to produce compatible spectra. We see that using the geometric solid angle significantly underestimates the true flux and, in fact, XMM-Newton should provide somewhat weaker or comparable constraints to eROSITA once this has been addressed.
@@
For annihilation, the revised XMM-Newton bounds are roughly an order of magnitude weaker than those from eROSITA across the full DM mass range. For decaying DM, XMM-Newton constraints are more competitive at low masses, where FSR dominates, but become nearly two orders of magnitude weaker at higher masses, where IC emission governs the size of the signal. For PBHs, we find that XMM-Newton constraints are consistently about an order of magnitude weaker than those from eROSITA over the full PBH mass range considered. This clarification realigns the role of XMM-Newton in the broader landscape of X-ray DM searches and confirms that eROSITA currently provides the strongest diffuse Galactic X-ray limits in the sub‑10 MeV regime, though other probes such as Voyager and 511 keV line studies remain more constraining in certain regions of parameter space.
@@ 6 Discussion and conclusions
We find that, even with just six months of data, eROSITA provides competitive X-ray constraints. For decaying DM below $\sim 10$ MeV, our limits surpass those from CMB and diffuse $\gamma$-ray observations, though they remain weaker than bounds from Voyager and Leo‑T dwarf galaxy heating in the same mass range. However, in this low-mass regime, signal predictions are subject to larger uncertainties due to charged particle propagation modeling. Looking ahead, eROSITA’s ability to resolve point sources will improve with future data releases, allowing for better subtraction of unresolved backgrounds and potentially yielding significantly stronger constraints. Current diffuse emission is dominated by unresolved sources such as coronal stellar emission and accreting white dwarfs. Improved estimates of this unresolved component could tighten the bounds presented here by orders of magnitude.
@@
In summary, eROSITA provides the strongest diffuse X-ray limits from annihilating and decaying dark matter in the $\lesssim 10$ MeV mass range, surpassing previous X-ray bounds from INTEGRAL, NuSTAR, and Suzaku. For decaying dark matter, our constraints also improve upon those from the CMB and diffuse $\gamma$-ray observations, though they remain weaker than the most stringent bounds from Voyager-1 and Leo-T dwarf galaxy heating. For PBHs, eROSITA limits are weaker than those from Voyager and 511 keV studies. Nevertheless, its diffuse all-sky coverage makes it uniquely suited to provide complementary and robust checks, especially with future deeper data releases.